Calculate de broglie wavelength and velocity of an electron carrying kinetic energy 2keV


1
Expert's answer
2021-01-30T16:47:44-0500

From the definition of the de Broglie wavelength, we have:


"\lambda=\dfrac{h}{p}=\dfrac{h}{mv}."

Let's find the velocity of the electron from the definition of the kinetic energy:


"KE=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE}{m}}."

Substituting "v" into the first equation, we get:


"\lambda=\dfrac{h}{\sqrt{2mKE}},"

"\lambda=\dfrac{6.626\cdot10^{-34}\ J\cdot s}{\sqrt{2\cdot9.1\cdot10^{-31}\ kg\cdot2\cdot10^3\ eV\cdot1.6\cdot10^{-19}\ \dfrac{J}{eV}}}=2.74\cdot10^{-11}\ m."

Finally, we can find the velocity of the electron:


"v=\sqrt{\dfrac{2\cdot2\cdot10^3\ eV\cdot1.6\cdot10^{-19}\ \dfrac{J}{eV}}{9.1\cdot10^{-31}\ kg}}=2.65\cdot10^7\ \dfrac{m}{s}."